OpenStudy (lopus):

to melt a snowball, with initial radius of 12 cm, its radius decreases at a constant rate of 0.5 cm / hr. if it starts to melt when t = 0, the rate of decline in the volume of the snowball after 12 hours is:

OpenStudy (anonymous):

The relation between radius and the volume is a cubic relation. If the radius is reducing at 0.5 cm/hr, the volume must decline at a constant rate of (0.5cm/hr)^3

OpenStudy (lopus):

\[volume=\frac{ 2 }{ 3 }\pi R^{3} \] \[\frac{ 2 }{ 3 }\pi (0.5*t)^{3}\] final t= 12

OpenStudy (lopus):

\[144\pi \frac{cm ^{3} }{hr }\] is possible?

OpenStudy (anonymous):

V = ( 4π/3 ) r³

OpenStudy (lopus):

the answer is : A. 50pi cm3/hr B. 60 pi .. C. 72 pi .. D. 120pi .. E. 144pi ..

OpenStudy (anonymous):

V= 4pi/3 * r^3 dV/dt = 4pi *r^2 *dr/dt dV/dt = 4pi*(6)^2 *.5 =72pi

OpenStudy (lopus):

6?